### v1

# bibfile.bib

Abstract
Let $\mathbb{F}$ be a non-archimedean local field, and let $G=\mathrm{SL}_{n}(\mathbb{F})$, $n\geq 3$. Let $X$ be an infinite-dimensional simply connected symmetric space of finite rank, with non-positive curvature operator. We prove that every continuous action by isometries of $G$ on $X$ has a fixed point.
1 Introduction
The main goal of this paper is to study continuous actions by isometries of the group $G=\mathrm{SL}_{n}(\mathbb{F})$, with $\mathbb{F}$ a non-archimedean local field, on simply connected infinite-dimensional symmetric spaces $X$ of non-positive curvature and finite rank, where the rank is defined as the dimension of a maximal embedded Euclidean subspace. Such spaces come from Hilbert spaces with indefinite bilinear forms, whose study dates back to Pontrjagin [pontrjagin], Iohvidov-Krein [iohkrein], and Naimark [naimark1, naimark2, naimark3, naimark4]. The study of the associated symmetric spaces was first suggested by Gromov ([gromov], Section 6). The particular case of the infinite-dimensional real hyperbolic space has been treated by Burger, Iozzi and Monod in [bim], where they construct equivariant embedding of trees into this space. More recently, infinite-dimensional symmetric spaces of finite rank have been widely studied by Duchesne [ducphd, duc2013, duc2015, duc2023].
In another paper [trees], we have studied actions of tree automorphism groups on such spaces $X$, showing that they can be irreducible only if $X$ has rank one. Here, we focus on actions of higher-rank algebraic groups, such as $G=\mathrm{SL}_{n}(\mathbb{F})$ with $n\geq 3$, on the same spaces $X$.
Our main goal will be to prove that, if $n\geq 3$, the group $G$ does not have any continuous fixed point-free action by isometries on $X$.
This is in contrast to the case $n=2$, where there exist continuous fixed point-free actions of $\mathrm{SL}_{2}(\mathbb{F})$ by isometries on the infinite-dimensional real hyperbolic space ([bim], Theorem C).
We start by precisely defining the family $\mathcal{X}$ of possible target spaces.
Definition 1.1.
A space $X\in\mathcal{X}$ is either a classical Riemannian manifold or a separable Hilbert manifold. It is a simply connected symmetric space (infinite-dimensional, in case it is a Hilbert manifold), it has non-positive curvature operator, and there is a finite bound on the dimension of isometrically embedded Euclidean spaces in $X$ (i. e. $X$ has finite rank).
Remark 1.2.
We have to include finite-dimensional symmetric spaces in the family $\mathcal{X}$, even though they are not the main focus of our paper, because we will prove our result by an induction that also relies on the finite-dimensional case.
Remark 1.3.
We can further say that, in the case where the space $X$ is finite-dimensional, the problem is already well understood. Indeed, let $f:G\rightarrow\text{Isom}(X)$ be any continuous homomorphism. Since $\text{Isom}(X)$ is a Lie group it has no small subgroups, and since $G$ is totally disconnected, by Van Dantzig’s theorem there is a basis of neighborhoods of the identity that are open compact subgroups. This implies that there exists an open compact subgroup of $G$ whose image in $\text{Isom}(X)$ is trivial. As $\mathrm{PSL}_{n}(\mathbb{F})$ is simple ([grove], Theorem 1.13), and the kernel of the projection of $\mathrm{SL}_{n}(\mathbb{F})$ on $\mathrm{PSL}_{n}(\mathbb{F})$ is finite, this implies that the homomorphism $f$ is trivial on a finite-index subgroup of $G$, i. e. the homomorphism has finite image. Then, since $X$ is a CAT(0) space, the existence of a bounded orbit implies the existence of a fixed point ([bh], II.2.8).
The first result in the infinite-dimensional case is due to Burger, Iozzi and Monod ([bim], Theorem C). In that paper the authors construct a family of equivariant embeddings of a regular tree $T_{d}$ (a tree where every vertex has degree $d\geq 3$) into an infinite-dimensional hyperbolic space, which corresponds to a family of continuous irreducible representations of the group $\mathrm{Aut}(T_{d})$ to $O(1,\infty)$, i. e. continuous irreducible representations to a separable Hilbert space that preserve a non-degenerate bilinear form of index $1$. This implies in particular that for every non-archimedean local field $\mathbb{F}$ the group $\mathrm{SL}_{2}(\mathbb{F})$, which is a subgroup of $\mathrm{Aut}(T_{q+1})$ where $q$ is the cardinality of the residue field of $\mathbb{F}$, has continuous fixed point-free actions by isometries on the infinite-dimensional real hyperbolic space.
In this paper, we focus on the higher-rank case, namely $G=\mathrm{SL}_{n}(\mathbb{F})$ with $n\geq 3$. We will show that the behavior seen in [bim] for $\mathrm{SL}_{2}(\mathbb{F})$ cannot happen in this case. Arturo Sánchez Gonzáles obtained a first result in this direction in his Ph.D. thesis, where he proved that for $n\geq 3$ there are no continuous irreducible representations of $\mathrm{SL}_{n}(\mathbb{Q}_{p})$ to $\mathrm{O}(2,\infty)$ ([arturo], Teorema C).
Our main result is the following:
Theorem 1.4.
Let $\mathbb{F}$ be a non-archimedean local field, and let $n\geq 3$. Let $G=\mathrm{SL}_{n}(\mathbb{F})$. Let $X\in\mathcal{X}$, as defined in Definition 1.1. Then any continuous action of $G$ by isometries on $X$ fixes a point in $X$.
Remark 1.5.
We can already observe that, if we can show the theorem for $n=3$, then the general case follows. In fact, it is shown in [cornulier] (Theorem 1.4, Proposition 1.2) that if $G$ is a simple linear algebraic group over a local field, then every action by isometries of $G$ on any metric space $X$ is either bounded, which means that the orbits are bounded, or metrically proper, which means that for every bounded subset $B\subset X$ the set $\{g\in G:gB\cap B\neq\emptyset\}$ has compact closure. We apply the result to $G=\mathrm{SL}_{n}(\mathbb{F})$. In case the action is bounded, we can already deduce that it has a fixed point in $X$, since our target spaces are CAT(0) ([bh], II.2.8). In case it is metrically proper, we take a closed subgroup $H<G$ isomorphic to $\mathrm{SL}_{3}(\mathbb{F})$. If $H$ has a fixed point $x\in X$, we know that the set $\{g\in G:gx=x\}$ has compact closure, which is a contradiction since it contains $H$ which is closed and non-compact.
Remark 1.6.
It is enough to show that the action of a cocompact subgroup $\Gamma<G$ is bounded. We will take $\Gamma$ to be a torsion-free cocompact lattice of $G$, that is, a torsion-free discrete cocompact subgroup of $G$. The existence of a discrete cocompact subgroup $\Gamma^{\prime}<G$ was proved in [borelharder] (Theorem A, Theorem 3.3), and due to ([garland], Theorem 2.7), there exists $\Gamma\lhd\Gamma^{\prime}$ of finite index that is also torsion-free.
Remark 1.7.
Remark 1.5 allows to extend Theorem 1.4 to a more general class of groups $G$, namely the class of all simple linear algebraic groups over a local field that contain $\mathrm{SL}_{3}(\mathbb{F})$ as a closed subgroup.
We conclude our introduction with a short outline of the paper.
After all the above reductions, our goal is to prove Theorem 1.4 for a torsion-free cocompact lattice $\Gamma$ in $G=\mathrm{SL}_{3}(\mathbb{F})$. First, we will show (Section 2) that every continuous action of $\Gamma$ by isometries on $X\in\mathcal{X}$ has a fixed point in $\overline{X}=X\cup\partial X$, where $\partial X$ is the boundary at infinity of $X$. This is done by generalizing a result that was proved by Wang [wang] in the finite-dimensional case. Finally, we will show (Section 3) that if a group $\Gamma$ with Kazhdan’s Property (T) has the property that whenever it acts continuously by isometries on a space $X\in\mathcal{X}$ it fixes a point in $\overline{X}$, and all its finite-index subgroups have the same property, then the group actually has to fix a point in $X$. This will allow us to conclude our proof.
This paper is part of my Ph.D. project under the supervision of Prof. Nicolas Monod at EPFL (Lausanne, Switzerland). I am grateful to Nicolas for proposing this project, for giving me valuable advice and feedback, and for sharing with me deep insights of the theory. I am also grateful to Bruno Duchesne and Pierre Py for their useful comments and suggestions on this paper.
2 Fixing a point in $\overline{X}$
In this section, we prove the following:
Theorem 2.1.
Let $\mathbb{F}$ be a non-archimedean local field, let $G=\mathrm{SL}_{3}(\mathbb{F})$, and let $\Gamma$ be a discrete, torsion-free, cocompact subgroup of $G$. Let $X\in\mathcal{X}$. Then every continuous action of $\Gamma$ by isometries on $X$ fixes a point in $\overline{X}$.
This result was proved by Wang in [wang] in the case where $X$ is a finite-dimensional Riemannian manifold of non-positive sectional curvature. The proof of our Theorem 2.1 will closely follow the proof in [wang], with some adaptations: most importantly, we need to take into account the fact that our space $X\in\mathcal{X}$ might also be infinite-dimensional.
2.1 The building of $G$ and the eigenvalue property
First, we introduce the Bruhat-Tits building $\Sigma$ of the group $G$, which is a Euclidean building of dimension two on which $G$ acts by isometries. It is a union of infinitely many Euclidean planes (apartments), each of them being tessellated by equilateral triangles (chambers). The action of $G$ is transitive in the set of pairs $(\mathcal{A},\mathcal{C})$ where $\mathcal{A}$ is an apartment and $\mathcal{C}$ is a chamber contained in $\mathcal{A}$. More details on the construction and structure of the building are explained in [buildings] (6.9).
As $\Sigma$ is a two-dimensional simplicial complex, for $i=0,1,2$ we may call $\Sigma(i)$ the set of $i$-simplices in $\Sigma$. The action of $G$ on $\Sigma(0)$ has three orbits, and each $2$-simplex $\sigma\in\Sigma(2)$ has exactly one vertex in each orbit.
For every vertex $v\in\Sigma(0)$, we can define its link graph $Lk(v)$. It is a graph whose vertices are the vertices in $\Sigma(0)$ adjacent to $v$ (i. e. connected to $v$ by an edge in $\Sigma(1)$), and where two vertices $u,w$ are connected by an edge if and only if there is a chamber in $\Sigma(2)$ with vertices $v,u,w$. We call $Lk(v)(0),Lk(v)(1)$ the set of vertices and edges of $Lk(v)$, respectively.
We describe the structure of the graph $Lk(v)$. As mentioned in [buildings] (6.9), it is the flag complex of $\mathrm{SL}_{3}(k)$, where $k$ is a finite field: the residue field of $\mathbb{F}$. This means that the vertices are points and lines in the projective plane over $k$, and whenever a point is contained in a line, the corresponding vertices are connected by an edge. If $q$ is the cardinality of the field $k$ (a power of a prime number), there are $N=2(q^{2}+q+1)$ vertices in total: $q^{2}+q+1$ points of $\mathbb{P}^{2}(k)$ and $q^{2}+q+1$ lines. Each vertex is connected to $q+1$ other vertices, thus there are in total $M=(q+1)(q^{2}+q+1)$ edges.
We show a property of the graph $Lk(v)$ that corresponds to the first eigenvalue property in [wang]. Let $v\in\Sigma(0)$ and let $Lk(v)$ be its link graph. After choosing an orientation for every edge in $Lk(v)(1)$, for every $f:Lk(v)(0)\rightarrow\mathbb{R}$ we can define $df:Lk(v)(1)\rightarrow\mathbb{R}$ by $df((u,w))=f(u)-f(w)$. We let $||f||^{2}=\sum_{u\in Lk(v)(0)}|f(u)|^{2}$ and $||df||^{2}=\sum_{(u,w)\in Lk(v)(1)}|f((u,w))|^{2}$.
Lemma 2.2.
There exists $\lambda>\frac{q+1}{2}$ such that, if $f$ satisfies $\sum_{u\in Lk(v)(0)}{f(u)}=0$, then $||df||^{2}\geq\lambda||f||^{2}$.
Proof.
Let $A$ be the adjacency matrix of $Lk(v)$. More explicitly, it is a matrix with $M$ rows and $N$ columns where, after an ordering of the vertices and of the edges has been chosen, the coefficient $A_{ij}$ is $1$ if the $i$-th edge starts from the $j$-th vertex, $-1$ if the $i$-th edge terminates at the $j$-th vertex, and 00 in all other cases. Then we can consider the function $f$ as a vector in $\mathbb{R}^{N}$ and the function $df$ as a vector in $\mathbb{R}^{M}$, so that we have $df=Af$. Therefore we have
$||df||^{2}=\langle df,df\rangle=\langle Af,Af\rangle=\langle f,A^{T}Af\rangle,$
where we use the standard scalar product in $\mathbb{R}^{M}$ and in $\mathbb{R}^{N}$.
In order to bound $\langle A^{T}Af,f\rangle$ in terms of $\langle f,f\rangle=||f||^{2}$, we study the eigenvalues of the matrix $B=A^{T}A$. It is a $N\times N$ square matrix where the coefficient $B_{ij}$ is $q+1$ if $i=j$, $-1$ if $i\neq j$ and the $i$-th and $j$-th vertex are connected by an edge, 00 if $i\neq j$ and the $i$-th and $j$-th vertex are not connected. We now describe the eigenspaces of $B$ with their respective eigenvalues:
  • If $f$ is a constant vector, then $Bf=0$ and thus $f$ is an eigenvector of $B$ with eigenvalue 00. This makes a $1$-dimensional eigenspace.
  • If $f$ takes a constant value $c$ on all points $p$ of the projective plane $\mathbb{P}_{2}(k)$, and the value $-c$ on all lines $l$ in $\mathbb{P}_{2}(k)$, then (recall that the graph is bipartite with respect to points and lines) $Bf=2(q+1)f$ and thus $f$ is an eigenvector of $B$ with eigenvalue $2(q+1)$. This makes another $1$-dimensional eigenspace.
  • Let $f$ be defined arbitrarily on the set of points $p$, with sum zero on this set. Then define $f$ at each line $l$ as $f(l)=\frac{1}{\sqrt{q}}\sum_{p\in l}{f(p)}$. Observe that, for every point $p$, we have
    $\frac{1}{\sqrt{q}}\sum_{l\ni p}{f(l)}=\frac{1}{q}\sum_{l\ni p}{\sum_{p^{\prime}\in l}{f(p^{\prime})}}=\frac{1}{q}\left(qf(p)+\sum_{p^{\prime}\in\mathbb{P}_{2}(k)}{f(p^{\prime})}\right)=f(p),$
    since when we sum over all points in all lines passing through $p$ the point $p$ is summed $q+1$ times while all the other points are summed once. We can conclude that, for every vertex $u\in Lk(v)(0)$, the sum of $f$ at all neighbors of $u$ in the graph is equal to $\sqrt{q}f(u)$. This implies that $Bf=(q+1-\sqrt{q})f$. As $f$ was chosen arbitrarily with sum zero on the set of points $p\in\mathbb{P}_{2}(k)$, this makes a $(q^{2}+q)$-dimensional eigenspace with eigenvalue $q+1-\sqrt{q}$.
  • In the previous construction, after fixing $f$ arbitrarily with sum zero on the set of points $p$ in $\mathbb{P}_{2}(k)$, we might also choose to define it on the set of lines by setting $f(l)=-\frac{1}{\sqrt{q}}\sum_{p\in l}{f(p)}$ for every line $l$. In this case we obtain $f(p)=-\frac{1}{\sqrt{q}}\sum_{l\ni p}{f(l)}$ for every point $p$, hence a $(q^{2}+q)$-dimensional eigenspace with eigenvalue $q+1+\sqrt{q}$.
With the above description, we have decomposed $\mathbb{R}^{N}$ as orthogonal direct sum of four eigenspaces for the operator $B$. What we need to show for the lemma is that, whenever $f$ is a vector in $\mathbb{R}^{N}$ with sum of the coordinates equal to zero, we have $\langle Bf,f\rangle\geq\lambda||f||^{2}$ for some $\lambda>\frac{q+1}{2}$. We can pick $\lambda=q+1-\sqrt{q}$, which is greater than $\frac{q+1}{2}$ for $q\geq 2$. Since the sum of the coordinates is zero the vector $f$ decomposes as an orthogonal sum of vectors in the second, third and fourth eigenspace described above, whose eigenvalues are all greater than or equal to $\lambda$. Hence the result follows.
Remark 2.3.
The function $f$ in Lemma 2.2 takes values in $\mathbb{R}$, but the result also holds when it takes values in a general Hilbert space. To see this, it suffices to restrict to the finite-dimensional subspace spanned by the (finitely many) values of $f$, then apply the lemma component-wise.
2.2 Existence of harmonic functions
We now define the most important tool for the proof of Theorem 2.1: harmonic functions on $\Sigma(0)$, which are $\Gamma$-equivariant functions from $\Sigma(0)$ to $X$ that minimize an energy functional.
An approach with harmonic functions was also used in [duc2015sup], where Duchesne proved a superrigidity result for cocompact lattices in higher rank semisimple Lie groups acting on spaces in the family $\mathcal{X}$.
Recall that $\Gamma$ is a discrete, torsion-free, cocompact subgroup of $G$. Its action on $\Sigma$ is free and proper ([garland], Lemma 2.6). Therefore, $\Gamma$ is the fundamental group of the finite simplicial complex $\sigma=\Sigma/\Gamma$, and $\Gamma$-equivariant functions $f:\Sigma(0)\rightarrow X$ correspond to functions $f_{\sigma}:\sigma(0)\rightarrow X$, where $\sigma(0)$ is the set of vertices in $\sigma$. The correspondence is defined by choosing a fundamental domain $D$ of $\Gamma$ in $\Sigma$ and assigning to $f_{\sigma}$ the values that $f$ takes at the vertices in $D$. Since $f$ is $\Gamma$-equivariant, all its values are uniquely determined by those at the vertices in $D$ and hence by $f_{\sigma}$.
Definition 2.4.
(Energy functional) The energy functional of a $\Gamma$-equivariant function $f:\Sigma(0)\rightarrow X$ is defined by
$E(f):=\sum_{(u,w)\in\sigma(1)}{d(f(\widetilde{u}),f(\widetilde{w}))^{2}},$
where $\sigma(1)$ is the set of edges in the $1$-skeleton of $\sigma$, and $\widetilde{u},\widetilde{w}\in\Sigma(0)$ are vertices connected by an edge in $\Sigma(1)$ that project to $u,w\in\sigma(0)$. Since $f$ is $\Gamma$-equivariant and the action of $\Gamma$ on $X$ is by isometries, $d(f(\widetilde{u}),f(\widetilde{w}))$ does not depend on the choice of $\widetilde{u},\widetilde{w}$.
Definition 2.5.
(Harmonic function) We say that the function $f$ is harmonic if it minimizes the energy functional among all $\Gamma$-equivariant functions $\Sigma(0)\rightarrow X$.
The first step of our proof of Theorem 2.1 is the following, which we prove in the rest of this subsection:
Proposition 2.6.
If the action of $\Gamma$ does not fix any point in the visual boundary $\partial X$, then there exists at least one harmonic function $f:\Sigma(0)\rightarrow X$.
Since a $\Gamma$-equivariant function $f$ is determined by the associated function $f_{\sigma}:\sigma(0)\rightarrow X$, we can consider the energy functional as a continuous function $E:X^{n}\rightarrow\mathbb{R}$, where $n$ is the cardinality of $\sigma(0)$. The function $E$ is bounded below by zero, and Proposition 2.6 is equivalent to $E$ having a minimum. Therefore, if $I\geq 0$ is the infimum of $E$ in $X^{n}$, we need to show that $E(x)=I$ for some $x\in X^{n}$.
As the product of CAT(0) spaces is still CAT(0), so is the space $X^{n}$. Moreover, since the space $X$ has finite telescopic dimension and this property is preserved by taking products, the space $X^{n}$ has finite telescopic dimension. We refer to [caply] for more information and results on finite telescopic dimension for CAT(0) spaces.
For all $m\in\mathbb{N}$, let $S_{m}:=\{x\in X^{n}:\ E(x)\leq I+1/m\}$. Since the energy functional is convex (the distance function in a CAT(0) space is convex), the sets $(S_{m})_{m\in\mathbb{N}}$ are a descending chain of closed convex subsets of $X^{n}$. It is proved in [caply] (Theorem 1.1) that in this case either the intersection $\bigcap_{m\in\mathbb{N}}{S_{m}}$ is non-empty or the intersection of the visual boundaries $\bigcap_{m\in\mathbb{N}}{\partial S_{m}}$ is a non-empty subset of $\partial X$.
If the intersection $\bigcap_{m\in\mathbb{N}}{S_{m}}$ is non-empty, any $x$ in this intersection must satisfy $E(x)=I$ and hence we are done. In the remaining case, let $\xi\in\partial X$ be in the intersection $\bigcap_{m\in\mathbb{N}}{\partial S_{m}}$. To complete the proof of Proposition 2.6, we show:
Proposition 2.7.
The point at infinity $\xi$ is fixed by the action of $\Gamma$.
Proof.
We can choose a sequence $(x_{m})_{m\in\mathbb{N}}$ of points in $X^{n}$ such that each $x_{m}$ is in $S_{m}$ and the sequence converges to $\xi$. For all $m\geq 1$, we have $E(x_{m})\leq I+1$.
Each point $x_{m}$ is associated with a $\Gamma$-equivariant function $f_{m}:\Sigma(0)\rightarrow X$. From the definition of the energy functional it follows that, whenever two vertices $v,w\in\Sigma(0)$ are connected by an edge in $\Sigma(1)$, we have $d(f_{m}(v),f_{m}(w))^{2}\leq E(f_{m})\leq I+1$.
Fix $g\in\Gamma$ and let $v_{1},\ldots,v_{n}$ be the vertices in the fundamental domain $D$. For every $m$, the point $x_{m}\in X^{n}$ has coordinates $(f_{m}(v_{1}),\ldots,f_{m}(v_{n}))$. From what we said above, for $1\leq i\leq n$ we have
$d(gf_{m}(v_{i}),f_{m}(v_{i}))=d(f_{m}(gv_{i}),f_{m}(v_{i}))\leq d_{\Sigma}(gv_{i},v_{i})\sqrt{I+1},$
where $d_{\Sigma}$ denotes the simplicial distance in $\Sigma$ (i. e. the graph distance in the $1$-skeleton). It follows that
$d(gx_{m},x_{m})\leq n\max_{1\leq i\leq n}{d_{\Sigma}(gv_{i},v_{i})}\sqrt{I+1}.$
This means that the distance $d(gx_{m},x_{m})$ is bounded above by a constant that does not depend on $m$. Therefore, since the sequence $(x_{m})_{m\in\mathbb{N}}$ converges to the point at infinity $\xi$, the sequence $(gx_{m})_{m\in\mathbb{N}}$ must also converge to $\xi$. This shows that $\xi$ is fixed by $g$.
2.3 Harmonic functions are constant
Our next goal is to show the following:
Proposition 2.8.
Let $f:\Sigma(0)\rightarrow X$ be harmonic. Then $f$ is constant.
Together with Proposition 2.6, this will prove Theorem 2.1: given a continuous action of $\Gamma$ by isometries on $X$, if it does not fix any point in $\partial X$ then Proposition 2.6 applies and we have a harmonic function $f:\Sigma(0)\rightarrow X$ that must be constant by Proposition 2.8, which means that the whole of $\Sigma(0)$ is sent to a single point $x\in X$ and therefore $x$ is fixed by $\Gamma$.
Therefore, we are left to prove Proposition 2.8. The proof is basically the same as in [wang]; we include it here in a form which is coherent with our situation and notation.
Definition 2.9.
(Differential) The differential of $f$ at $v\in\Sigma(0)$ is defined to be
$Df|_{v}:Lk(v)(0)\rightarrow T_{f(v)}X$
such that $\mathrm{exp}(Df|_{v}(u))=f(u)$ for all $u\in Lk(v)(0)$, where $\mathrm{exp}$ is the exponential map on $T_{f(v)}X$.
Recall that a $\Gamma$-equivariant function $f:\Sigma(0)\rightarrow X$ is uniquely determined by its values at the vertices $v_{1},\ldots,v_{n}$ in the fundamental domain $D$. In order for the map to be harmonic, the variation of the energy caused by any infinitesimal variation of $f(v_{i})$, $1\leq i\leq n$, must be zero. This is expressed in the following lemma:
Lemma 2.10.
Let $f$ be harmonic. Then
$\sum_{u\in Lk(v_{i})(0)}{Df|_{v_{i}}(u)}=0$
for $1\leq i\leq n$.
Proof.
We consider the variation at $v_{i}$ given by $y\in T_{f(v_{i})}X$. We have:
$\frac{d}{dt}E(t)$ $=\frac{d}{dt}\left(\sum_{u\in Lk(v_{i})(0)}{d(f(v_{i}),f(u))^{2}}\right)=$
$=\frac{d}{dt}\left(\sum_{u\in Lk(v_{i})(0)}{d(\mathrm{exp}(ty),f(u))^{2}}\right)=$
$=\sum_{u\in Lk(v_{i})(0)}{\langle y,-2Df|_{v_{i}}(u)\rangle},$
where the last equality holds because $f(v_{i})$ and $f(u)$ are connected by a minimal geodesic in $X$ and the gradient of the function $x\mapsto d(x,u)^{2}$ at $x=v_{i}$ is $-2Df|_{v_{i}}(u)$.
Since $f$ is harmonic it is a critical point of the energy functional, hence
$\sum_{u\in Lk(v_{i})(0)}{\langle y,Df|_{v_{i}}(u)\rangle}=0.$
Since $y$ was arbitrary in $T_{f(v_{i})}X$, the lemma follows.
We can now proceed with the proof of Proposition 2.8. Let $f$ be harmonic. Let $i\in\{1,\ldots,n\}$ and let $u,w\in Lk(v_{i})(0)$. Since $X$ is non-positively curved, we have
$||Df|_{v_{i}}(u)-Df|_{v_{i}}(w)||\leq d(f(u),f(w)).$
Therefore,
$\sum_{(u,w)\in Lk(v_{i})(1)}{||Df|_{v_{i}}(u)-Df|_{v_{i}}(w)||^{2}}\leq\sum_{(u,w)\in Lk(v_{i})(1)}{d(f(u),f(w))^{2}}.$
Denoting the left-hand side by $||d(Df)_{v_{i}}||^{2}$ and summing over $i$, we obtain
$\sum_{i=1}^{n}{||d(Df)_{v_{i}}||^{2}}\leq\sum_{i=1}^{n}{\sum_{(u,w)\in Lk(v_{i})(1)}{d(f(u),f(w))^{2}}}.$ (1)
We claim that the right-hand side is equal to $(q+1)E(f)$. Using the notation of Definition 2.4, we see that for every edge $(u,w)\in\sigma(1)$ we are summing $d(f(\widetilde{u}),f(\widetilde{w}))^{2}$ once for every vertex $v\in\sigma(0)$ such that $u,w,v$ are vertices of a simplex in $\sigma(2)$. The number of such vertices is always $q+1$, since it is the same as the degree of the vertex $w$ in the link graph $Lk(u)$. Therefore, (1) becomes:
$\sum_{i=1}^{n}{||d(Df)_{v_{i}}||^{2}}\leq(q+1)E(f).$ (2)
Consider now, for $1\leq i\leq n$, the function $g_{i}=Df|_{v_{i}}$. It is a function $g_{i}:Lk(v_{i})(0)\rightarrow T_{f(v_{i})}X$, and we know by Lemma 2.10 that $\sum_{u\in Lk(v_{i})(0)}{g_{i}(u)}=0$. We can apply Remark 2.3 and get $||dg_{i}||^{2}\geq\lambda||g_{i}||^{2}$ for some $\lambda>\frac{q+1}{2}$. This means that
$\sum_{(u,w)\in Lk(v_{i})(1)}{||Df|_{v_{i}}(u)-Df|_{v_{i}}(w)||^{2}}\geq\lambda\sum_{u\in Lk(v_{i})(0)}{||Df|_{v_{i}}(u)||^{2}}.$
Since $||Df|_{v_{i}}(u)||^{2}=d(f(v_{i}),f(u))^{2}$, we see that if we sum over $i$, we get $2E(f)$ on the right-hand side, since for each edge $(v,u)\in\sigma(1)$ we are summing the quantity $d(f(\widetilde{v}),f(\widetilde{u}))^{2}$ twice. Therefore, we have:
$\sum_{i=1}^{n}{||d(Df)_{v_{i}}||^{2}}\geq 2\lambda E(f).$ (3)
From (2) and (3) we get $(q+1)E(f)\geq 2\lambda E(f)$, which, since $\lambda>\frac{q+1}{2}$ and $E(f)\geq 0$, implies $E(f)=0$. By the definition of $E(f)$, this implies that $f$ is constant on $\Sigma(0)$, proving Proposition 2.8. As explained above, Theorem 2.1 follows.
3 Obtaining a fixed point inside $X$
In this section, we complete the proof of Theorem 1.4. We know from Theorem 2.1 that the action of the discrete torsion-free cocompact subgroup $\Gamma<G=\mathrm{SL}_{3}(\mathbb{F})$ on any $X\in\mathcal{X}$ has to fix a point in $\overline{X}$. For convenience, we name this property as in [wang]:
Definition 3.1.
(Property (F)) A group $\Gamma$ is said to have Property (F) if any continuous action by isometries of $\Gamma$ on any space $X\in\mathcal{X}$ has a fixed point in $\overline{X}$.
We also use the following definition:
Definition 3.2.
(Property (FH)) A locally compact topological group $\Gamma$ is said to have Property (FH) if any continuous affine isometric action of $\Gamma$ on a Hilbert space has a fixed point.
Remark 3.3.
If $\Gamma$ is $\sigma$-compact, Property (FH) is equivalent to Kazhdan’s Property (T). This is known as the Delorme-Guichardet Theorem ([kazhdanT], Theorem 2.12.4).
In this section, we show the following general result, which will allow us to conclude the proof of Theorem 1.4:
Theorem 3.4.
Let $\Gamma$ be a locally compact group. Assume that every finite-index subgroup of $\Gamma$ has Property (F) and Property (FH). Then any continuous action by isometries of $\Gamma$ on a space $X\in\mathcal{X}$ has a fixed point in $X$.
3.1 More about the family $\mathcal{X}$
As a first step, we determine more explicitly the family of possible spaces $X$. On one hand, it contains all simply connected finite-dimensional symmetric spaces of non-positive curvature. On the other hand, it is shown in [duc2015] (Corollary 1.10) that any space in $\mathcal{X}$, after possibly removing the Euclidean de Rham factor, is isometric to a finite product of irreducible (finite-dimensional) symmetric spaces of non-compact type and spaces of the form $\mathrm{O}(p,\infty)/\mathrm{O}(p)\times\mathrm{O}(\infty)$, $\mathrm{U}(p,\infty)/\mathrm{U}(p)\times\mathrm{U}(\infty)$, $\mathrm{Sp}(p,\infty)/\mathrm{Sp}(p)\times\mathrm{Sp}(\infty)$ with $p\geq 1$, up to homothety. Such spaces were studied in [duc2013]. We use the notation $X_{p}(\mathbb{K})$ for these spaces, where $\mathbb{K}$ is the field of real, complex, or quaternionic numbers.
A general $X\in\mathcal{X}$ is then a product
$X\cong\mathbb{R}^{r}\times\prod_{i=1}^{h}{M_{i}}\times\prod_{i=1}^{k}{N_{i}}$ (1)
where each $M_{i}$ is a simply connected finite-dimensional symmetric space with non-positive curvature and each $N_{i}$ is homothetic to some $X_{p}(\mathbb{K})$.
We can reduce our problem to each of these single factors. According to a result by Foertsch and Lytchak [foly], every geodesic space $X$ of finite affine rank has a unique decomposition
$X=Y_{0}\times\prod_{i=1}^{n}{Y_{i}}$
where $Y_{0}$ is a Euclidean space (possibly a point) and each $Y_{i}$, $1\leq i\leq n$, is an irreducible metric space not isometric to the real line, nor to a point. Moreover, each isometry of $X$ is a product of isometries of the single $Y_{i}$, up to possibly permuting factors which are isometric. This result was inspired by the classical De Rham decomposition for finite-dimensional Riemannian manifolds [derham].
We will prove in the Appendix that the spaces $X_{p}(\mathbb{K})$ are irreducible metric spaces. Since they are infinite-dimensional, they cannot appear as factors of the finite-dimensional symmetric spaces ($\mathbb{R}^{r}$ or $M_{i}$) in the decomposition (1) of $X$. Therefore, the result of Foertsch and Lytchak implies that any isometry of $X$ is a product of an isometry of $\mathbb{R}^{r}\times\prod_{i=0}^{h}{M_{i}}$ and an isometry of each factor $N_{i}$, $1\leq i\leq k$, up to possibly permuting isometric factors. The classical result of De Rham then allows us to conclude that any isometry is a product of isometries of each individual factor in the decomposition, up to possibly permuting isometric factors.
There is also an alternative way to see that the factors that appear in the decomposition (1) of $X$ coincide with the factors of Foertsch-Lytchak and therefore any isometry of $X$ is a product of isometries of the individual factors up to possibly permuting isometric ones. Indeed, since $X$ is a geodesic space and all its geodesics extend indefinitely in both directions, the same must be true for the factors $(Y_{i})_{0\leq i\leq n}$ in the Foertsch-Lytchak decomposition (this follows from the fact that geodesics in a product of metric spaces project to geodesics in factors, as seen in [bh], I.5.3). Then, as a consequence of Proposition 2.2 in [duc2023], these factors must be totally geodesic submanifolds of $X$ and therefore coincide (up to the order) with the factors in the decomposition (1). This was suggested by Bruno Duchesne.
Going back to the statement we wish to prove, Theorem 3.4, we see that $\Gamma$ must have a finite-index subgroup $\Gamma^{\prime}$ that acts on $X$ without permuting the factors. If we prove that the action of $\Gamma^{\prime}$ has a fixed point in $X$, then the action of $\Gamma$ will have a finite orbit and hence a fixed point ([bh], II.2.8). Therefore, from now on, we may just assume that the action of $\Gamma$ does not permute the factors.
Now, $\Gamma$ acts on each factor and the action on $X$ is just the product of these actions. The action on the Euclidean factor has a fixed point due to Property (FH). Each $M_{i}$ and each $N_{i}$ is a symmetric space that belongs to the class $\mathcal{X}$, hence the action of $\Gamma$ has a fixed point either on the space or on its boundary (Property (F)). If we prove Theorem 3.4 in the case where $X$ is a single factor, we will find that in the general case the action on each factor has a fixed point, and therefore the global action has a fixed point, proving the theorem. Hence from now on we will assume that $X$ is either an irreducible finite-dimensional symmetric space of non-compact type, or a space $X_{p}(\mathbb{K})$ for some $p\geq 1$ and $\mathbb{K}\in\{\mathbb{R},\mathbb{C},\mathbb{H}\}$.
3.2 Property (FH) and soluble normal subgroups
The main tool to prove Theorem 3.4 will be a form of Levi decomposition for parabolic subgroups of $H=\text{Isom}(X)$, where a parabolic subgroup will be expressed as a semidirect product of a soluble normal subgroup and a Levi factor of lower rank. In this subsection, we prove a lemma that will allow us to restrict the action to the Levi factor, and thus to reduce the rank of the target space.
Lemma 3.5.
Let $\Gamma$ be a locally compact group with Property (FH), and let $H=N\rtimes_{\phi}Q$ be a semidirect product of topological groups. Assume that $N$ is soluble and that every Abelian quotient that arises from its derived series is isomorphic (as a topological group) to the additive group of a Hilbert space, where the action of $Q$ (induced from $\phi$) preserves a scalar product. Then any continuous homomorphism $f:\Gamma\rightarrow H$ has image contained in a conjugate of $Q$.
Proof.
We prove the lemma by induction on the solubility length of $N$. The base case will be the trivial case with solubility length zero, where $N$ is the trivial group, and $H=Q$. For the inductive step, we assume that $N$ has solubility length $n\geq 1$ and that the lemma is true in all cases with solubility length $\leq n-1$.
Let $N^{\prime}:=[N,N]$ be the commutator subgroup of $N$, which has solubility length $n-1$, let $A=N/N^{\prime}$ be the abelianization of $A$, and let $ab:N\rightarrow A$ be the natural projection (with kernel $N^{\prime}$). We recall that $Q$ acts on $N$ by the map $\phi:Q\rightarrow\text{Aut}(N)$ that appears in the semidirect product and that associates to every $q\in Q$ the conjugation by $q$. Since the commutator subgroup $N^{\prime}$ is characteristic, it is preserved by the action of $Q$, which means that the action descends to the quotient $A$. We call $\phi_{ab}$ this induced action. By hypothesis, $A$ is isomorphic (as a topological group) to the additive group of a Hilbert space, and the action $\phi_{ab}$ of $Q$ preserves a scalar product in this Hilbert space.
We can now define an action of $\Gamma$ on $A$ as follows: for every $g\in\Gamma$ there exist unique $n(g)\in N,q(g)\in Q$ such that $f(g)=n(g)q(g)$; then we set
$F(g)(a)=ab(n(g))\phi_{ab}(q(g))(a)\ \ \ \forall a\in A.$
The action of $Q$ is by automorphisms, so by linear operators, while the action of $N$ is by translations. Therefore, the action of $\Gamma$ is by affine transformations, and as the action of $Q$ preserves a scalar product, the whole action of $\Gamma$ is by isometries with respect to the metric induced by this scalar product on $A$.
Using Property (FH), we find that the action of $\Gamma$ on $A$ fixes a point $a_{0}$, which means that if we conjugate $f$ by an element $n_{0}\in N$ with $ab(n_{0})=a_{0}$ we get an action that fixes the neutral element of $A$, which means that a conjugate of the homomorphism $f$ has image contained in $N^{\prime}\rtimes Q$. Now, since $N^{\prime}$ has solubility length $n-1$, by inductive hypothesis this conjugate has image contained in a conjugate of $Q$, so in the end we find that $f$ also has image contained in a conjugate of $Q$.
3.3 The induction in the finite-dimensional case
The proof of Theorem 3.4 will be by induction on the rank of the target space $X$. The base case (rank 0) is trivial: a simply connected symmetric space of rank 0 is reduced to a point because otherwise it would contain a geodesic that is an embedded copy of $\mathbb{R}$. Therefore, it suffices to prove our theorem when $X$ has rank $p\geq 1$ assuming that it is true for spaces of rank $\leq p-1$.
As the group $\Gamma$ has Property (F), we know that every action of $\Gamma$ on $X$ has to fix a point either in $X$ or in $\partial X$. The first case is the thesis of our theorem, so from now on we will assume that we are in the second case, i. e. $\Gamma$ fixes $x\in\partial X$. This means that there is a homomorphism $f:\Gamma\rightarrow H_{x}$, where $H_{x}<H=\text{Isom}(X)$ is the parabolic subgroup defined by $H_{x}=\{h\in H:hx=x\}$.
In this subsection, we focus on the case where $X$ is finite-dimensional. We will use the classical theory of Lie groups and symmetric spaces of non-compact type, for which a good reference is the book by Eberlein [eberlein]. We use results from (2.17) of that book, which we report here using our notation.
At the beginning, Eberlein defines a parabolic subgroup $H_{x}$ of $H$ as the stabilizer of a point at infinity $x$, then defines a homomorphism $T_{x}:H_{x}\rightarrow H$ by fixing a point $p\in X$ and a tangent vector $V$ at $p$ pointing to $x$, and setting
$T_{x}(h)=\lim_{t\rightarrow+\infty}{e^{-tV}he^{tV}}.$
Then he proves (2.17.5) that there is a decomposition
$H_{x}=N_{x}Z_{x}$
where $N_{x}$ is the kernel of the homomorphism $T_{x}$ and $Z_{x}=\{h\in H:he^{tV}=e^{tV}h\ \forall t\in\mathbb{R}\}\subseteq H_{x}$. The subgroup $N_{x}$ is connected and normal in $H_{x}$, and the decomposition of any $h\in H_{x}$ as the product of an element in $N_{x}$ and an element in $Z_{x}$ is unique. It follows that $H_{x}$ is a semidirect product
$H_{x}=N_{x}\rtimes Z_{x}.$
Eberlein then defines $\mathfrak{n}_{x}$ as the Lie algebra of $N_{x}$ and shows that it is nilpotent (2.17.14), which means that the group $N_{x}$ is nilpotent in the algebraic sense ([hilneeb], Theorem 11.2.5). Moreover, as shown in the proof of the Iwasawa decomposition ([hilneeb], Theorem 13.3.8), $N_{x}$ is simply connected.
Now, if we have the homomorphism $f:\Gamma\rightarrow H_{x}$, we can compose it with the projection $\pi:H_{x}\rightarrow Z_{x}$ and get a homomorphism $\overline{f}:\Gamma\rightarrow Z_{x}$. In the proof of ([eberlein], 2.17.5) it is shown that if $\gamma_{px}$ is a geodesic that passes through $p$ and points to $x$, the set $F(\gamma_{px})$ defined as the union of all geodesics parallel to $\gamma_{px}$ is preserved by the action of $Z_{x}$. Therefore, we have an action of $\Gamma$ on $F(\gamma_{px})$. It is shown in [eberlein] (2.11.4) that $F(\gamma_{px})$ is a complete totally geodesic submanifold of $X$ and
$F(\gamma_{px})=\mathbb{R}^{r}\times F_{S}(\gamma_{px})$
where $r\geq 1$ and $F_{S}(\gamma_{px})$ is a symmetric space of non-compact type.
As $\Gamma$ acts on $F(\gamma_{px})$, the result of De Rham [derham] implies that the action splits as the product of an action on $\mathbb{R}^{r}$ and an action on $F_{S}(\gamma_{px})$. Due to Property (FH), the action on $\mathbb{R}^{r}$ fixes a point; hence there exists an embedded copy of $F_{S}(\gamma_{px})$ that is preserved by $\Gamma$ and on which $\Gamma$ acts.
As $\Gamma$ has Property (F) it must fix a point either in $F_{S}(\gamma_{px})$ or in its boundary. Since $F(\gamma_{px})$ is embedded in $X$ and since $r\geq 1$, the rank of $F_{S}(\gamma_{px})$ must be strictly lower than the rank of $X$, so we can use our inductive hypothesis and find that $\Gamma$ has to fix a point inside $F_{S}(\gamma_{px})$.
This means that the image of the homomorphism $\overline{f}$ is contained in a compact subgroup $Q$ of $Z_{x}$, and therefore the image of $f$ is contained in $N_{x}\rtimes Q$. Now we would like to apply Lemma 3.5, which would allow us to conclude that the image of $f$ is actually contained in a conjugate of $Q$, and therefore that the action of $\Gamma$ fixes a point in $X$, which is the thesis of Theorem 3.4.
To verify the hypotheses of Lemma 3.5, we need to see that each Abelian quotient that appears in the derived series of the soluble group $N_{x}$ is isomorphic to the additive group of a Hilbert space where the action of $Q$ preserves a scalar product. We know from [hilneeb] (Proposition 11.2.4) that the groups in the derived series of a connected Lie group $\mathcal{G}$ are integral subgroups (that is, they are generated by a subalgebra of the Lie algebra of $\mathcal{G}$), and that integral subgroups $\mathcal{H}$ of a simply connected soluble Lie group $\mathcal{G}$ are closed and simply connected, with $\mathcal{G}/\mathcal{H}$ isomorphic to $\mathbb{R}^{\dim{\mathcal{G}/\mathcal{H}}}$ ([hilneeb], Proposition 11.2.15). We can apply this to the Lie group $\mathcal{G}=N_{x}$, finding that the groups in the derived series are closed and simply connected, and each quotient is isomorphic to $\mathbb{R}^{n}$ for some $n$. We also know that the groups in the derived series are characteristic subgroups of $N_{x}$ (meaning that they are preserved by any automorphism of $N_{x}$), and hence the action of $Q$ given by the semidirect product $N_{x}\rtimes Q$ descends to an action on each quotient. Furthermore, since $Q$ is compact, each quotient has a scalar product that is preserved by this action of $Q$: it suffices to start from any scalar product and integrate it over $Q$ to obtain a $Q$-invariant one. This completes the proof.
3.4 The infinite-dimensional case
We are left with showing the inductive step for the infinite-dimensional case, where we have $X=X_{p}(\mathbb{K})$ for some $p\geq 1$ and $\mathbb{K}\in\{\mathbb{R},\mathbb{C},\mathbb{H}\}$. Since $\Gamma$ acts by isometries on $X$, it has a homomorphism to the group $H=\mathrm{PO}_{\mathbb{K}}(p,\infty)$, where $\mathrm{PO}_{\mathbb{R}}(p,\infty)=\mathrm{PO}(p,\infty)$, $\mathrm{PO}_{\mathbb{C}}(p,\infty)=\mathrm{PU}(p,\infty)$, and $\mathrm{PO}_{\mathbb{H}}(p,\infty)=\mathrm{PSp}(p,\infty)$. In fact, it is shown in [duc2023] (Theorem 3.3) that $\mathrm{Isom}(X_{p}(\mathbb{K}))=\mathrm{PO}_{\mathbb{K}}(p,\infty)$, except in the case $\mathbb{K}=\mathbb{C}$ where $\mathrm{PU}(p,\infty)$ is a subgroup of index two in $\mathrm{Isom}(X_{p}(\mathbb{C}))$. In that case, we have a subgroup $\Gamma^{\prime}<\Gamma$ of index $\leq 2$ with a homomorphism to $\mathrm{PU}(p,\infty)$, and if we show that $\Gamma^{\prime}$ fixes a point in $X$, then this implies that $\Gamma$ has a finite orbit and hence fixes a point.
Therefore, we consider a continuous homomorphism $f:\Gamma\rightarrow H=\mathrm{PO}_{\mathbb{K}}(p,\infty)$, that is, a continuous action of $\Gamma$ by projective transformations on a $\mathbb{K}$-vector space $V$ of countable dimension that preserves a nondegenerate sesquilinear form of index $p$.
Since $\Gamma$ has Property (F), we can assume that the action on $X=X_{p}(\mathbb{K})$ associated with $f$ fixes a point $x\in\partial X$, therefore, the image of $f$ is contained in $H_{x}=\{h\in H:hx=x\}$.
In the course of the proof, we will consider matrices associated to elements of $\mathrm{PO}_{\mathbb{K}}(p,\infty)$ or other projective groups. This means that we are writing the explicit matrix form of a representative of the element in the corresponding matrix group (in this case, $\mathrm{O}_{\mathbb{K}}(p,\infty)$). Moreover, in the case $\mathbb{K}=\mathbb{H}$, we use the convention for which the multiplication by a scalar in the vector space is on the right, so that the action is on the left.
We can choose a basis $(e_{1},e_{2},\ldots)$ of $V$ such that the sesquilinear form is expressed by the infinite matrix
$\large\left(\begin{array}[]{c|c|c}0&Id_{p}&0\\ \hline\cr Id_{p}&0&0\\ \hline\cr 0&0&Id_{\infty}\end{array}\right)$
where $Id_{p}$ is the identity $p\times p$ matrix and $Id_{\infty}$ is the identity infinite matrix.
Each point at infinity of $X_{p}(\mathbb{K})$ is associated with a flag of isotropic subspaces of the vector space $V$: this was shown in [duc2013] (Proposition 6.1) in the real case, but is true in all cases, as observed in [duc2023] (as a consequence of Remark 3.1 in [duc2023]).
We consider the flag associated with the point at infinity $x$ (which is fixed by the action of $\Gamma$). Let $q\leq p$ be the dimension of the maximal subspace in the flag. We may choose the basis in such a way that this maximal subspace is exactly $\text{Span}\{e_{1},\ldots,e_{q}\}$. Then we can change the order of the elements $(e_{q+1},e_{q+2},\ldots)$ of the basis in such a way that the sesquilinear form is now expressed by the matrix
$\large\left(\begin{array}[]{c|c|c}0&Id_{q}&0\\ \hline\cr Id_{q}&0&0\\ \hline\cr 0&0&\Large J\end{array}\right),\ \ \ J\normalsize=\left(\begin{array}[]{c|c}-Id_{p-q}&0\\ \hline\cr 0&Id_{\infty}\end{array}\right).$
Now we consider the matrix associated with a transformation in $H_{x}$ with respect to this new basis, and find some restrictions on its structure. We see it as a block matrix similar to the one expressing the sesquilinear form, i. e. with two blocks of dimension $q$ and one block of infinite dimension, both horizontally and vertically. This means that we are splitting the vector space $V$ into a direct sum $V_{1}\oplus V_{2}\oplus V_{3}$ where $V_{1}$ and $V_{2}$ have dimension $q$ and $V_{3}$ has infinite dimension, and each block of the matrix corresponds to a continuous linear map from $V_{i}$ to $V_{j}$ for some $i,j\in\{1,2,3\}$.
First of all, we note that to preserve $V_{1}$ the two bottom blocks in the left column must vanish. For the upper left block, we generally have a block-upper triangular structure depending on the flag associated with $x$, but we are not going to need that structure in our proof, so we may just assume that it belongs to some subgroup $S<\mathrm{GL}_{q}(\mathbb{K})$ (possibly the whole of $\mathrm{GL}_{q}(\mathbb{K})$). We get the form:
$\large\left(\begin{array}[]{c|c|c}M&Y&Z\\ \hline\cr 0&L&W\\ \hline\cr 0&B&R\end{array}\right).$
In order to preserve the sesquilinear form we must have
$\left(\begin{array}[]{c|c|c}M^{T}&0&0\\ \hline\cr Y^{T}&L^{T}&B^{T}\\ \hline\cr Z^{T}&W^{T}&R^{T}\end{array}\right)\left(\begin{array}[]{c|c|c}0&Id_{q}&0\\ \hline\cr Id_{q}&0&0\\ \hline\cr 0&0&J\end{array}\right)\left(\begin{array}[]{c|c|c}M&Y&Z\\ \hline\cr 0&L&W\\ \hline\cr 0&B&R\end{array}\right)=\left(\begin{array}[]{c|c|c}0&Id_{q}&0\\ \hline\cr Id_{q}&0&0\\ \hline\cr 0&0&J\end{array}\right)$
which is equivalent to
$\left(\begin{array}[]{c|c|c}0&M^{T}L&M^{T}W\\ \hline\cr L^{T}M&L^{T}Y+Y^{T}L+B^{T}JB&L^{T}Z+Y^{T}W+B^{T}JR\\ \hline\cr W^{T}M&W^{T}Y+Z^{T}L+R^{T}JB&W^{T}Z+Z^{T}W+R^{T}JR\end{array}\right)=\left(\begin{array}[]{c|c|c}0&Id_{q}&0\\ \hline\cr Id_{q}&0&0\\ \hline\cr 0&0&J\end{array}\right).$
From the first row we immediately get $L=(M^{T})^{-1}$ and $W=0$, so we might simplify to
$\left(\begin{array}[]{c|c|c}0&Id_{q}&0\\ \hline\cr Id_{q}&M^{-1}Y+(M^{-1}Y)^{T}+B^{T}JB&M^{-1}Z+B^{T}JR\\ \hline\cr 0&(M^{-1}Z)^{T}+R^{T}JB&R^{T}JR\end{array}\right)=\left(\begin{array}[]{c|c|c}0&Id_{q}&0\\ \hline\cr Id_{q}&0&0\\ \hline\cr 0&0&J\end{array}\right).$
From $R^{T}JR=J$ we see that $R\in\mathrm{O}_{\mathbb{K}}(p-q,\infty)$. From $M^{-1}Z+B^{T}JR=0$ we get $Z=-MB^{T}JR$. Therefore, the matrix of our transformation has the form
$\large\left(\begin{array}[]{c|c|c}M&Y&-MB^{T}JR\\ \hline\cr 0&(M^{T})^{-1}&0\\ \hline\cr 0&B&R\end{array}\right)$
with $R\in\mathrm{O}_{\mathbb{K}}(p-q,\infty)$ and $Y$ satisfying $M^{-1}Y+(M^{-1}Y)^{T}=-B^{T}JB$.
Our next step will be to decompose the group $H_{x}$ as a semidirect product, using the explicit matrix form. We can project $H_{x}$ to the (projective) group of block-diagonal matrices, with block dimensions $q,q,\infty$, using the map
$\left(\begin{array}[]{c|c|c}M&Y&-MB^{T}JR\\ \hline\cr 0&(M^{T})^{-1}&0\\ \hline\cr 0&B&R\end{array}\right)\mapsto\left(\begin{array}[]{c|c|c}M&0&0\\ \hline\cr 0&(M^{T})^{-1}&0\\ \hline\cr 0&0&R\end{array}\right).$
It is immediate to see that this is a well-defined group homomorphism. As $M\in S<GL_{q}(\mathbb{K})$ and $R\in\mathrm{O}_{\mathbb{K}}(p-q,\infty)$, this gives a projection $H_{x}\rightarrow\mathrm{P}(S\times\mathrm{O}_{\mathbb{K}}(p-q,\infty))$.
The kernel $N$ of this projection is the space of matrices having the form
$\large\left(\begin{array}[]{c|c|c}Id_{q}&Y&-B^{T}J\\ \hline\cr 0&Id_{q}&0\\ \hline\cr 0&B&Id_{\infty}\end{array}\right)$
with $Y+Y^{T}=-B^{T}JB$. We will show that this group is soluble with solubility length two and that it satisfies the assumptions of Lemma 3.5.
The map sending a matrix as above to the block $B$ determines a homomorphism $N\rightarrow A$ where $A$ is the additive group of $q\times\infty$ matrices (i. e. linear maps from the $q$-dimensional vector space $V_{2}$ to the infinite-dimensional Hilbert space $V_{3}$), which is Abelian and isomorphic with the direct sum of $q$ Hilbert spaces of countable dimension, which is isomorphic with a single Hilbert space of countable dimension. The kernel of the map is the set of matrices in $N$ with $B=0$, which is isomorphic with the additive group of $q\times q$ matrices $Y$ that satisfy $Y+Y^{T}=0$, i. e. the additive group of skew-symmetric matrices, isomorphic with $\mathbb{K}^{q(q-1)/2}$.
Therefore, we have a semidirect product decomposition
$H_{x}=N\rtimes\mathrm{P}(S\times\mathrm{O}_{\mathbb{K}}(p-q,\infty))$
where $N$ is soluble.
What we need to show for our theorem is that any continuous homomorphism $f:\Gamma\rightarrow H_{x}$ has image contained in the stabilizer of a point in the symmetric space $X$. Consider the homomorphisms $f_{S}:\Gamma\rightarrow\mathrm{P}S$ and $f_{O}:\Gamma\rightarrow\mathrm{PO}_{\mathbb{K}}(p-q,\infty)$ defined by $f_{S}=\pi_{S}\circ f$ and $f_{O}=\pi_{O}\circ f$, where $\pi_{S}$ and $\pi_{O}$ are the natural projections of $H_{x}$ onto $\mathrm{P}S$ and $\mathrm{PO}_{\mathbb{K}}(p-q,\infty)$ respectively.
The homomorphism $f_{S}$ has image in $\mathrm{P}S<\mathrm{PGL}_{q}(\mathbb{K})$. As $\mathrm{PGL}_{q}(\mathbb{K})$ is a Lie group of rank $q-1<p$, we can use the inductive hypothesis of Theorem 3.4 and see that the image of $f_{S}$ fixes a point in the associated symmetric space. Up to conjugating the homomorphism (which is the same as changing the basis of $V_{1}$ and $V_{2}$ that we consider in our matrix decomposition), we may assume that the image of $f_{S}$ is contained in $\mathrm{PO}_{\mathbb{K}}(q)$.
Similarly, we see that the image of $f_{O}$ is contained in a subgroup that fixes a point in the infinite-dimensional symmetric space $X_{\mathbb{K}}(p-q,\infty)$. Up to conjugating, we may therefore assume that the image of $f_{O}$ is contained in $\mathrm{P}(\mathrm{O}_{\mathbb{K}}(p-q)\times\mathrm{O}_{\mathbb{K}}(\infty))$, meaning that every element is block-diagonal for the decomposition of $V_{3}$ used in the definition of the matrix $J$, and that the two nonzero blocks that appear are in $\mathrm{O}_{\mathbb{K}}(p-q)$ and $\mathrm{O}_{\mathbb{K}}(\infty)$.
Now we know that the image of $f$ is contained in $N\rtimes Q$ with $Q=\mathrm{P}(\mathrm{O}_{\mathbb{K}}(q)\times\mathrm{O}_{\mathbb{K}}(p-q)\times\mathrm{O}_{\mathbb{K}}(\infty))$. We wish to apply Lemma 3.5 to find that it is actually contained in a conjugate of $Q$, which would prove our theorem. We only need to check the hypothesis that the conjugation of elements in $N$ by elements of $Q$ preserves a Hilbert scalar product in the two Abelian groups that appear in the soluble group structure, namely the additive group $A$ of maps $V_{2}\rightarrow V_{3}$ (the block $B$ in the decomposition) and the additive group of skew-symmetric matrices $Y$.
We show that the standard scalar product in these spaces is preserved. The norm associated to this scalar product can be regarded, in both cases, as the Hilbert sum of the squares of all coefficients of the associated matrix (in the case of skew-symmetric matrices, summing over the whole matrix instead of just on the upper-triangular part multiplies all the norms by $2$). The conjugation of a generic element in $N$ by a generic element in $S\times\mathrm{O}_{\mathbb{K}}(p-q,\infty)$ gives
$\left(\begin{array}[]{c|c|c}M&0&0\\ \hline\cr 0&(M^{T})^{-1}&0\\ \hline\cr 0&0&R\end{array}\right)$ $\left(\begin{array}[]{c|c|c}Id_{q}&Y&-B^{T}J\\ \hline\cr 0&Id_{q}&0\\ \hline\cr 0&B&Id_{\infty}\end{array}\right)\left(\begin{array}[]{c|c|c}M^{-1}&0&0\\ \hline\cr 0&M^{T}&0\\ \hline\cr 0&0&R^{-1}\end{array}\right)=$
$=\left(\begin{array}[]{c|c|c}Id_{q}&MYM^{T}&-MB^{T}JR^{-1}\\ \hline\cr 0&Id_{q}&0\\ \hline\cr 0&RBM^{T}&Id_{\infty}\end{array}\right).$
In our case $M\in O_{\mathbb{K}}(q)$, and this also implies $M^{T}\in O_{\mathbb{K}}(q)$. When a matrix $Y$ is multiplied by $M^{T}$ on the right, every row is transformed with an orthogonal transformation and therefore its norm stays the same. Hence, the norm of the whole matrix remains the same. Similarly, when multiplying by $M$ on the left, every column is transformed orthogonally and the general norm remains the same.
As for the transformation $B\mapsto RBM^{T}$, the same remark applies for the right multiplication by $M^{T}$, and also for the left multiplication by $R$ since $R\in\mathrm{O}_{\mathbb{K}}(p-q)\times\mathrm{O}_{\mathbb{K}}(\infty)$ and hence it preserves the standard scalar product on $V_{3}$. This concludes the proof of Theorem 3.4.
3.5 Conclusion
We are now ready to prove Theorem 1.4.
Let $G=\mathrm{SL}_{3}(\mathbb{F})$, and let $\Gamma$ be a discrete, torsion-free, cocompact subgroup of $G$, which exists by ([borelharder], Theorem A, Theorem 3.3) and ([garland], Theorem 2.7). By Theorem 2.1, $\Gamma$ has Property (F). Since finite-index subgroups of $\Gamma$ are again discrete, torsion-free, and cocompact in $G$, they also have Property (F).
The group $G$ has Kazhdan’s Property (T) ([kazhdanT], 1.4), and since this property is inherited by lattices ([kazhdanT], Theorem 1.7.1), the subgroup $\Gamma$ and all its finite-index subgroups also have Kazhdan’s Property (T). By the Delorme-Guichardet Theorem ([kazhdanT], Theorem 2.12.4), this means that these groups have Property (FH).
By Theorem 3.4, any continuous action by isometries of $\Gamma$ on a space $X\in\mathcal{X}$ has a fixed point in $X$. As explained in Remark 1.6, this implies that any continuous action by isometries of $G$ on $X$ has a fixed point, proving Theorem 1.4.
Appendix A Appendix
In this appendix we prove that the spaces $X_{p}(\mathbb{K})$, as defined at the beginning of Section 3.1, are irreducible metric spaces, meaning that they are not isometric to the product of two metric spaces not reduced to a point.
Set $X=X_{p}(\mathbb{K})$. We suppose by contradiction that $X=X_{1}\times X_{2}$ for some metric spaces $X_{1},X_{2}$ not reduced to a point. First, we note that $X_{1},X_{2}$ must be geodesic spaces, since $X$ is a geodesic space and geodesics in a product project to geodesics in factors ([bh], I.5.3). Moreover, geodesics in $X_{1},X_{2}$ must extend indefinitely, since geodesics in $X$ extend indefinitely. This implies that the boundary at infinity of $X_{1}$ and $X_{2}$ is non-empty. We also note that $X_{1},X_{2}$ must be CAT(0) spaces, as they are convex subsets of the CAT(0) space $X$, and that they must be complete, as their product is complete.
From now on, we will assume that $X$ has rank at least $2$. In fact, since both $X_{1}$ and $X_{2}$ contain an infinite geodesic, $X$ must contain an isometrically embedded copy of $\mathbb{R}^{2}$.
It is proven in [bh] (II.9) that the Tits boundary $\partial_{T}(X_{1}\times X_{2})$ of the product of two complete CAT(0) spaces is the spherical join of the Tits boundaries of the factors. Therefore, it suffices to prove:
Proposition A.1.
For $X=X_{p}(\mathbb{K})$, the Tits boundary $\partial_{T}X$ is not the spherical join of two non-empty metric spaces.
The Tits boundary of $X_{p}(\mathbb{K})$ was studied by Duchesne [duc2013]: it is a thick spherical building, and all finite configurations of simplices are the same that appear in the spherical building of the finite-dimensional analogue $\mathrm{O}_{\mathbb{K}}(p,q)$ for sufficiently large $q$. Details on the structure of these buildings can be found in the book by Abramenko and Brown [buildings] (6.7). We show the following:
Lemma A.2.
In the spherical building of $\mathrm{O}_{\mathbb{K}}(p,q)$, every simplex has diameter strictly smaller than $\pi/2$.
Proof.
We refer to [buildings] for details on the construction of the building. Here we only describe the structure of the fundamental apartment, which allows us to see that every simplex has diameter $<\pi/2$.
[…]
We now describe the structure of the simplices in the sphere, which is forced by symmetry. We can choose coordinates such that the sphere is the unit sphere in $\mathbb{R}^{p}$, for $1\leq i\leq p$ the subspace $\mathrm{Span}\{e_{i}\}$ is associated with the point with $i$-th coordinate equal to $1$, and the subspace $\mathrm{Span}\{f_{i}\}$ is associated with the point with $i$-th coordinate equal to $-1$. These points determine a subdivision of the sphere into $2^{p}$ super-simplices: each of them is the intersection of the sphere with the subset of $\mathbb{R}^{p}$ resulting from a specific choice of the signs of the $p$ coordinates. Each super-simplex corresponds to a maximal isotropic subspace, and is then subdivided barycentrically into $p!$ individual simplices. Each subset $A\subset\{e_{1},\ldots,e_{p},f_{1},\ldots,f_{p}\}$ with $\mathrm{Span}(A)$ totally isotropic is associated with the barycenter of the set of points associated with elements of $A$.
Clearly, all super-simplices are isometric, and all individual simplices are isometric. Let $\mathcal{S}$ be a super-simplex, without loss of generality,
$\mathcal{S}=\{(x_{1},\ldots,x_{p})\in S^{p-1}:x_{i}\geq 0,1\leq i\leq p\}.$
We now describe the individual simplices that make up $\mathcal{S}$. For each permutation $\sigma$ of $\{1,\ldots,p\}$, we have the simplex
$s_{\sigma}=\{(x_{1},\ldots,x_{p})\in\mathcal{S}:x_{\sigma(1)}\geq x_{\sigma(2)}\geq\cdots\geq x_{\sigma(p)}\}.$
This corresponds to the flag
$\mathrm{Span}\{e_{\sigma(1)}\}\subset\mathrm{Span}\{e_{\sigma(1)},e_{\sigma(2)}\}\subset\cdots\subset\mathrm{Span}\{e_{\sigma(1)},\ldots,e_{\sigma(p)}\}.$
We assume without loss of generality $\sigma=id$ and show that the simplex $s_{id}$ has diameter $<\pi/2$. Let $x\in S^{p-1}$. The set of points in $S^{p-1}$ at distance $\geq\pi/2$ from $x$ can be defined by $S^{p-1}\cap\{y\in\mathbb{R}^{p}:\langle x,y\rangle\leq 0\}$, where we use the standard scalar product in $\mathbb{R}^{p}$. If we assume that there exist $x,y\in s_{id}$ with $d(x,y)\geq\pi/2$, they must satisfy $\langle x,y\rangle\leq 0$. As all their coordinates are non-negative and their first coordinate is at least $\sqrt{1/p}$, we have $\langle x,y\rangle\geq 1/p$, a contradiction. Hence, the diameter of $s_{id}$, and similarly the diameter of any simplex in the sphere, is strictly smaller than $\pi/2$.
Now we can deduce Proposition A.1 from the following general result:
Theorem A.3.
Let $\Sigma$ be a thick spherical building in which every simplex has diameter strictly smaller than $\pi/2$. Then $\Sigma$ is not the spherical join of two non-empty metric spaces.
Proof.
Suppose by contradiction that $\Sigma$ is the spherical join $A_{1}*A_{2}$. Then $A_{1},A_{2}$ are non-empty subsets of $\Sigma$ with the following properties (deduced from the construction of the spherical join, see [bh], I.5.13):
  1. (a)
    for all $x\in A_{1},y\in A_{2}$ we have $d(x,y)=\pi/2$;
  2. (b)
    for all $z\in\Sigma$ there exist $x\in A_{1},y\in A_{2}$ such that $z$ belongs to the unique geodesic segment between $x$ and $y$; if $z\not\in A_{1}\cup A_{2}$, these $x,y$ are the unique points in $A_{1},A_{2}$ at minimal distance from $z$.
Moreover, since $\Sigma$ is a CAT(1) space ([bh], II.10A.4), $A_{1}$ and $A_{2}$ must also be CAT(1) spaces ([bh], II.3.15). In particular they are $\pi$-geodesic.
Let $x\in A_{1},y\in A_{2}$. By (a) they are at distance $\pi/2$, thus there exists a unique geodesic segment between them. Since $\Sigma$ is a spherical building, this segment has to be part of a geodesic loop $\gamma$ of length $2\pi$. Let $x^{\prime}$ be the point on this loop that is antipodal to $x$. Let $z_{1}\in\gamma$ such that $d(y,z_{1})=\pi/2-\varepsilon$ and $d(z_{1},x^{\prime})=\varepsilon$, where $\varepsilon$ is small enough so that the geodesic from $y$ to $z_{1}$ extends in a unique way at least up to $x^{\prime}$. Let $x_{1}\in A_{1},y_{1}\in A_{2}$ be the points obtained by applying property (b) on $z_{1}$. We have $d(x,y_{1})=\pi/2$ by property (a), hence $d(y_{1},z_{1})\geq|d(x,z_{1})-d(x,y_{1})|=\pi/2-\varepsilon$ by the triangle inequality. Since $y_{1}$ is the unique point in $A_{2}$ at minimal distance from $z_{1}$, and since $d(y,z_{1})=\pi/2-\varepsilon$, we must have $y_{1}=y$, and $x_{1}=x^{\prime}$ since it must be at distance $\epsilon$ from $z_{1}$ on a geodesic extending the one coming from $y$ and the geodesic $\gamma$ is the unique extension at least up to the point $x^{\prime}$. Therefore $x^{\prime}\in A_{1}$.
Now, since simplices in $\Sigma$ have diameter strictly smaller than $\pi/2$, and since the building is thick, there exists a point $r$ in the interior of the geodesic segment between $y$ and $x^{\prime}$ where the geodesic ramifies, meaning that the extension beyond $r$ of the geodesic segment from $y$ to $r$ is not unique. The point $z_{1}$ is of course on the segment between $r$ and $x^{\prime}$. The non-uniqueness of the geodesic extension beyond $r$ means that there exists another geodesic $\widetilde{\gamma}$ that contains the points $x,y,r$ but not the point $x^{\prime}$. It will contain instead another point $x^{\prime\prime}$ antipodal to $x$. We can choose $z_{2}$ on the geodesic segment between $r$ and $x^{\prime\prime}$, close enough to $x^{\prime\prime}$ so that the geodesic segment from $r$ to $z_{2}$ extends uniquely at least up to $x^{\prime\prime}$, and repeat the same proof as before to show that $x^{\prime\prime}\in A_{1}$.
Now we have $d(x^{\prime},x^{\prime\prime})\leq d(x^{\prime},r)+d(r,x^{\prime\prime})=\pi-2d(y,r)<\pi$, hence since $A_{1}$ is $\pi$-geodesic the geodesic segment between $x^{\prime}$ and $x^{\prime\prime}$ is contained in $A_{1}$. Let $m$ be the midpoint of this segment. From the CAT(1) inequality applied on the triangle with vertices $r,x^{\prime},x^{\prime\prime}$ we get $d(r,m)<\pi/2-d(y,r)$. Therefore $d(y,m)\leq d(y,r)+d(r,m)<\pi/2$, which is a contradiction since $y\in A_{2}$ and $m\in A_{1}$.
[…]
[heading=bibintoc]